Ricoh is out and Konica-Minolta is in… Building wide, copiers will be changed out between now and January. M.E.R. #26 will be used for some temporary new machine storage and prep/staging during the transition. Busted by the popo this work will be completed elsewhere
John E Green pulled both pumps and removed large amounts of towels around the impellers. They reinstalled the pumps and the electricians verified amp draw of around 4 amps. Pumps have been put back into normal operation.
warmed and blewdown boilers ,switched #2 boiler as lead Ran chemical pumps for boilers in manual all day to bring up levels. blowdown is closed for #2 boiler (conductivity 1000 ) normal shift
Ordered new micro switch for the door, and installed two new screws in door plate.
Plant rounds and readings in the Energy Center. Warmed #1 Boiler. Blew down #3 Boiler and checked LWCO. Tested boiler water and cooling towers water. Med gas Manifold readings. Metasys Checks and Cold and/or warm calls. Added 1 gal. of Halochlor to Yard Tower. Warmed #2 Boiler put in Auto in the PWC. And Auto remote at boiler panel. Adjusted surface Blow heat exchanger. from 100 deg. to 90 deg. Worked on Health Stream. Filled Softener Brine tank after regeneration.